Use the following functions and answer the questions? f(x)=4x^2-9x+2
a. Does the parabola open up or down?
b. What is the vertex?
c. What is the axis of symmetry? x=
d. What are the x-intercept (include the factors that you used to find them)
e. What are the y-intercept (include the factors that you used to find them)

Answer by stanbon(75887)

Use the following functions and answer the questions? f(x)=4x^2-9x+2
a. Does the parabola open up or down?
Ans: Up because the 4x^2 term is positive.
-------------------------------------------------
b. What is the vertex?
Ans: Vertex occurs when x = -b/2a = 9/8
f(9/8) = -3.063
Vertex: (1.125,-3.063}
-------------------------------
c. What is the axis of symmetry? x = 1.125
-----------------------------------------------
d. What are the x-intercept (include the factors that you used to find them)
Let y = 0 and solve for "x":
4x^2 -9x + 2 = 0
(4x^2 -8x -x + 2) = 0
4x(x-2) -(x-2) = 0
(x-2)(4x-1) = 0
x = 2 or x = 1/4
===========================
e. What are the y-intercept (include the factors that you used to find them)
Let x = 0 and solve for "y":
f(x)=4x^2-9x+2
y = 0 + 0 + 2
y = 2
====================
Cheers,
Stan H.